A colony of bacteria originally contains 200 bacteria. It doubles in size every 30 minutes. How many hours will it take for the colony to contain 2,000 bacteria? (Round your answer to one decimal place.
Answer by Alan3354(69443)

A colony of bacteria originally contains 200 bacteria. It doubles in size every 30 minutes. How many hours will it take for the colony to contain 2,000 bacteria?
------------
t = # of 1/2 hour increments.
---
2000 = 200*2^t
2^t = 10
t*log(2) = log(10) = 1
t = 1/log(2) =~ 3.322 half hours.
etc